# Freeze-in and Freeze-out production of Higgs Portal Majorana Fermionic Dark Matter during and after Reheating

### v1

@@ Abstract
In this paper, we investigate the production of Majorana fermionic dark matter (DM) via the Higgs portal, considering both freeze-in and freeze-out mechanisms during and after the post-inflationary reheating phase. We assume that the Universe is reheated through the decay of the inflaton into a pair of massless fermions. Our analysis focuses on how the non-standard evolution of the Hubble expansion rate and the thermal bath temperature during reheating influence DM production. Additionally, we examine the impact of electroweak symmetry breaking (EWSB), distinguishing between scenarios where DM freeze-in or freeze-out occurs before or after EWSB. We further explore the viable DM parameter space and its compatibility with current and future detection experiments, including XENONnT, LUX-ZEPLIN (LZ), XLZD, and collider searches. Moreover, we incorporate constraints from the Lyman-$\alpha$ bound to ensure consistency with small-scale structure formation.
@@ I Introduction
One compelling alternative is the freeze-in mechanism McDonald (2002); Hall et al. (2010); Elahi et al. (2015); Bernal et al. (2017), where DM interacts so feebly with the SM bath that it never reaches thermal equilibrium. Instead, its relic abundance is generated gradually through the decay and annihilations of SM particles. Due to their extremely weak couplings, such DM candidates are often referred to as feebly interacting massive particles (FIMPs). Another simple way to evade the experimental constraints on WIMP is to consider non-standard cosmological histories, such as the post-inflationary reheating epoch. In recent years, the study of non-standard cosmological histories has gained substantial interest in DM phenomenology Maity and Saha (2018); Garcia et al. (2020, 2021a, 2021b); Giudice et al. (2001); Barman et al. (2022); Bhattiprolu et al. (2022); Harigaya et al. (2014, 2019); Okada and Seto (2021); Ghosh et al. (2022); Haque and Maity (2022); Ahmed et al. (2023); Bernal and Xu (2022); Bernal et al. (2024a, 2019a, 2019b); Bhatia and Mukhopadhyay (2021); Hamdan and Unwin (2018); Arcadi et al. (2024); Barman et al. (2024a); González et al. (2024a, b); Silva-Malpartida et al. (2025); Bernal et al. (2024b); Barman et al. (2024b); Cosme et al. (2024a, b); Haque et al. (2024a, b).
@@
In our work, we investigate the production of Majorana fermionic DM ($\chi$) via the Higgs portal interaction $\bar{\chi}\chi H^{\dagger}H/\Lambda$, where $H$ represents the SM Higgs doublet, and $\Lambda$ serves as the cutoff scale. While some recent studies Lebedev et al. (2024); Arcadi et al. (2024); Ikemoto et al. (2023); Haba et al. (2024); Biswas et al. (2020) have explored the impact of reheating on Majorana DM production, they often assume instantaneous reheating scenarios and treat the reheating temperature as the maximum bath temperature. However, in realistic inflationary models, reheating occurs over a finite duration, where the maximum temperature can be significantly higher than the reheating temperature Giudice et al. (2001). Since DM production may depend on both temperatures, our analysis adopts a non-instantaneous reheating scenario for a more accurate description of the process. Furthermore, we consider both freeze-in and freeze-out production mechanisms during the reheating phase. During this phase, the inflaton field ($\phi$) oscillates around the minima of its potential and gradually decays into SM particles, thereby reheating the Universe. We assume that the inflaton oscillates with a monomial potential, $V(\phi)\propto\phi^{2\,n}$, leading to an average equation of state (EoS) given by $w_{\rm\phi}=(n-1)/(n+1)$ Turner (1983). Depending on the choice of $n$ and the inflaton decay width, the energy densities of the inflaton and the radiation bath follow a non-trivial evolution during reheating. This modified cosmological history can significantly alter the DM production process, affecting both its relic abundance and detection prospects.
@@ II Reheating dynamics
After the end of inflation, the inflaton field $(\phi)$ starts to oscillate around the minima of its potential, and transfers its energy into massless fields known as radiation. The process of transferring energy from the inflaton field ($\phi$) to the radiation through perturbative decays is called reheating. Here, we assume that the inflaton potential during reheating takes the form $V(\phi)\propto\phi^{2\,n}$ with n being an integer, which could originate from $\alpha$-attractor inflation models Kallosh and Linde (2013); Kallosh et al. (2013)
@@
Here, $M_{\rm pl}(=2.4\times 10^{18}$ GeV) is the reduced Planck mass, $\Lambda$ is the mass-scale fixed by the CMB power spectrum, which is typically of the order $8\times 10^{15}$ GeV, and the parameter $(\alpha,n)$ controls the shape of the potential. The coherently oscillating inflaton field can be decomposed as Shtanov et al. (1995); Ichikawa et al. (2008)
@@
$\omega=m_{\phi}(t)\,\xi\,,~{}~{}~{}\mbox{where}~{}~{}\xi=\sqrt{\frac{\pi\,n}{2\,(2\,n-1)}}\frac{\Gamma\left(\frac{1}{2}+\frac{1}{2\,n}\right)}{\Gamma\left(\frac{1}{2\,n}\right)}\,.$ (2)
@@
$m_{\phi}^{\rm end}\simeq\sqrt{\frac{{4n\,\left(2n-1\right)}}{{3\,\alpha}}}\frac{\Lambda^{\frac{2}{n}}}{M_{\rm pl}}\left(\rho_{\phi}^{\rm end}\right)^{\frac{n-1}{2n}}\,.$ (4)
The average energy density of the inflaton ($\rho_{\rm\phi}$) is defined with respect to $\phi_{0}$ as $\rho_{\phi}=(1/2)\langle({\dot{\phi}}^{2}+V(\phi))\rangle=V(\phi)|_{\phi_{0}}$. During reheating, the inflaton transfers its energy density to SM radiation energy density $(\rho_{\rm R})$. In order to solve reheating dynamics, the Boltzmann equations for the inflaton and radiation energy density supplemented with the Hubble equation are,
@@
where, dots denote derivatives with respect to cosmic time $t$, and $\Gamma_{\phi}$ is the inflaton decay rate. We consider inflaton decays into a pair of massless fermions $f$ and $\bar{f}$ through the interaction $y\,\phi\,\bar{f}\,f$, where $y$ is the dimensionless Yukawa coupling. Incorporating the inflaton oscillation effect, the effective inflaton decay rate $\Gamma_{\phi}$ can be written as Garcia et al. (2021a); Haque et al. (2023); Silva-Malpartida et al. (2023); Bernal and Xu (2022),
@@
The numerical values of $\sum\nu^{3}\lvert\mathcal{P}_{\nu}\rvert^{2}$ for different values of $w_{\rm\phi}$ are provided in table-1. Incorporating all these parameters, we have numerically solved the coupled Boltzmann equation (Eq. 5) with the appropriate initial conditions. For the analytical solution, we assume the inflaton decay width remains negligible compared to the Hubble rate throughout the reheating phase. Under this assumption, the decay term on the right-hand side of Eq. 5a can be neglected, and a straightforward integration yields the following solution,
@@
where $H_{\rm end}$ denotes the Hubble constant at the end of the inflation. In thermal equilibrium, the temperature of the inflaton decay products is given by $\rho_{\rm R}=\epsilon\,T^{4}$, where $\epsilon=(\pi^{2}/30)\,g_{\star}$, with $g_{\star}$ representing the effective number of relativistic degrees of freedom in the thermal bath. For an analytical estimation of the temperature, we assume $g_{\star}\sim\mathcal{O}(100)$, which remains a reasonable approximation for temperatures below $1$ GeV. In our numerical analysis, we solve the following equation along with Eqs. 5, to obtain the correct evolution of the temperature,
@@
$T_{\rm max}$ $=$ $\left[\frac{6\,M_{\rm pl}^{2}\,(1+w_{\phi})m^{end}_{\phi}\,H_{end}}{8\,\pi\,\epsilon\,(5-9w_{\phi})}y_{\rm eff}^{2}\right]^{1/4}\left[\left(\frac{8}{3+9w_{\phi}}\right)^{\frac{3(1+3w_{\phi}}{9w_{\phi}-5}}-\left(\frac{8}{3+9w_{\phi}}\right)^{\frac{8}{9w_{\phi}-5}}\right]^{1/4}$ (12)
@@
To avoid disrupting Big Bang Nucleosynthesis (BBN), the reheating temperature must satisfy $T_{\rm re}\gtrsim T_{\rm BBN}=4$ MeV Kawasaki et al. (2000); Hannestad (2004); Barbieri et al. (2025); de Salas et al. (2015). The upper bound on the reheating temperature is set by the inflationary scale, with $T_{\rm re}\lesssim 10^{15}$ GeV, derived from the constraint on the Hubble parameter during inflation, $H_{\rm I}^{\rm CMB}\leq 2\times 10^{-5}M_{\rm pl}$ Akrami et al. (2020); Ade et al. (2021). However, the thermal bath temperature can be expressed in terms of $T_{\rm re}$
${}T(a)\simeq T_{\rm re}\left\{\begin{array}[]{ll}&\left(\frac{A}{A_{\rm re}}\right)^{-\frac{3}{8}(1+3w_{\rm\phi})}~{}~{}~{}\mbox{for}~{}~{}A_{\rm max}\leq a\leq A_{\rm re}\\ &\left(\frac{A}{A_{\rm re}}\right)^{-1}~{}~{}~{}\mbox{for}~{}~{}A\geq A_{\rm re}\,,\end{array}\right.$ (17)
@@
${}H(a)\simeq H_{\rm re}\left\{\begin{array}[]{ll}&\left(\frac{A}{A_{\rm re}}\right)^{-\frac{3}{2}(1+w_{\rm\phi})}~{}~{}~{}\mbox{for}~{}~{}A_{\rm max}\leq A\leq A_{\rm re}\\ &\left(\frac{A}{A_{\rm re}}\right)^{-2}~{}~{}~{}\mbox{for}~{}~{}A\geq A_{\rm re}\,,\end{array}\right.$ (20)
@@ III.1 DM production before EWSB
[…]
@@
${}\langle\sigma v\rangle=\frac{g^{2}}{n_{\text{eq}}^{2}}\cdot\frac{T^{6}}{32\,\pi^{5}\,\Lambda^{2}}\,I_{1}(m_{\chi},T)$ (26)
@@ III.1.1 Freeze-in
${}\frac{N_{\chi}(A)}{dA}=\frac{A^{2}}{H}\langle\sigma v\rangle(n_{\rm eq})^{2}.$ (27)
@@
${}N_{\rm{\chi}}(A)=\frac{g^{2}}{24\,(3-7\,w_{\rm\phi})\,\pi^{5}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T_{\rm re}^{6}}{H_{\rm re}}\left[\left(\frac{A}{A_{\text{re}}}\right)^{\frac{3(3-7w_{\phi})}{4}}-\left(\frac{A_{\rm max}}{A_{\text{re}}}\right)^{\frac{3(3-7w_{\phi})}{4}}\right]$ (29)
From the above solution, $03-7w_{\rm\phi}>$ 0 when $w_{\rm\phi}<3/7$, therefore most of the production occurs around $A\sim A\,(T\sim m_{\rm\chi})$, consequently, the freeze-in temperature $T_{\rm fi}\sim m_{\chi}$  [^1] This type of freeze-in scenario is known as infra-red (IR) freeze-in Bernal et al. (2017); Yaguna (2011); Chu et al. (2012); Blennow et al. (2014); Merle and Totzauer (2015); Shakya (2016); Hessler et al. (2017); König et al. (2016); Duch et al. (2018); Heeba et al. (2018); Lebedev and Toma (2019), where bulk of the DM production occurs at $T\sim m_{\rm\chi}$.. But, on the other hand, $03-7w_{\rm\phi}<$ 0 when $w_{\rm\phi}>3/7$, most of the production occurs around $A\sim A_{\rm max}$ and this leads to the freeze-in temperature $T_{\rm fi}\sim T_{\rm max}$ 222 DM production that occurs predominantly at very high temperatures, typically near the maximum temperature of the Universe, is referred to as Ultraviolet (UV) freeze-in Hall et al. (2010); Elahi et al. (2015); McDonald (2016); Chen and Kang (2018); Biswas et al. (2020); Bernal et al. (2019b, 2020a, 2020b); Barman et al. (2020, 2021).. These two scenarios are illustrated in Fig. 3 (dashed lines). In the left plot, for $w_{\rm\phi}=0$, the co-moving number density is frozen at $m_{\chi}/T\sim 1$, and in the right side, for $w_{\rm\phi}=0.5$, the co-moving number density is frozen at the beginning of reheating (at $T\sim T_{\rm max}$).
Figure 3: The evolution of the co-moving DM density for freeze-in as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$ (left), $w_{\rm\phi}=0.5$ (right) with $T_{\rm re}=10^{9}$ GeV. Each plot presents three different cases: $m_{\rm\chi}<T_{\rm re}$ (solid), $T_{\rm re}<m_{\rm\chi}<T_{\rm max}$ (dashed) and $m_{\rm\chi}>T_{\rm max}$ (dot-dashed). The vertical red line corresponds to $T\sim T_{\rm fi}$, and the vertical black line corresponds to $T=T_{\rm re}$.
@@
$N_{\rm{\chi}}(A)=\frac{g^{2}\,m_{\rm\chi}}{64\,(1+3\,w_{\rm\phi})\,\pi^{4}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T^{5}_{\text{re}}}{H_{\text{re}}}\left(\frac{T_{\text{re}}}{T_{\text{max}}}\right)^{\frac{7-11w_{\rm\phi}}{1+3w_{\rm\phi}}}e^{-\frac{2\,m_{\rm\chi}}{T_{\rm max}}}$ (30)
@@
$N_{\rm{\chi}}(A)=\frac{g^{2}}{32\,\pi^{5}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T^{6}_{\text{re}}}{H_{\text{re}}}\left(1-\frac{A_{\text{re}}}{A}\right)$ (31)
Although $m_{\rm\chi}<T_{\rm re}$, the DM production continues up to $T\sim T_{\rm re}$ point, then its production becomes negligible. This is true for $w_{\rm\phi}<3/7$ (see solid line in the left Fig. 3) and the freeze-in temperature $T_{\rm fi}\sim T_{\rm re}$. On the other hand, when $w_{\rm\phi}>3/7$, the DM production is frozen just at the beginning of reheating, which we discussed earlier (see solid line in the right Fig. 3) and the freeze-in temperature $T_{\rm fi}\sim T_{\rm max}$
@@
$\Omega_{\rm{\chi}}=\Omega_{R}h^{2}\frac{\sqrt{3}M_{\rm pl}\,g^{2}}{24\,\pi^{5}\epsilon^{\frac{3}{2}}\,T_{0}}\left\{\begin{array}[]{ll}&\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\left(\frac{m_{\rm\chi}}{T_{\rm re}}\right)^{\frac{14\,w_{\phi}-6}{1+3w_{\phi}}}\,,~{}~{}~{}\mbox{for}~{}~{}~{}T_{\rm max}>m_{\rm\chi}>T_{\rm re}\,,w_{\rm\phi}<3/7\\ &\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\,,~{}~{}~{}\mbox{for}~{}~{}~{}m_{\rm\chi}<T_{\rm re}\,,w_{\rm\phi}<3/7\\ &\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\left(\frac{T_{\rm max}}{T_{\rm re}}\right)^{\frac{(7\,w_{\phi}-3)}{4}}\,,~{}~{}~{}\mbox{for}~{}~{}~{}m_{\rm\chi}<T_{\rm max}\,,w_{\rm\phi}>3/7\\ &\frac{3\pi}{8}\frac{m_{\chi}^{2}}{(1+3w_{\phi})\Lambda^{2}}\left(\frac{T_{\text{re}}}{T_{\text{max}}}\right)^{\frac{7-11w_{\rm\phi}}{1+3w_{\rm\phi}}}e^{-\frac{2\,m_{\rm\chi}}{T_{\rm max}}}\,,~{}~{}~{}\mbox{for}~{}~{}m_{\rm\chi}>T_{\rm max}\,.\end{array}\right.$ (36)

### v2

@@ Abstract
In this paper, we investigate the production of Majorana fermionic dark matter (DM) via the Higgs portal, considering both freeze-in and freeze-out mechanisms during and after the post-inflationary reheating phase. We assume that the Universe is reheated through the decay of the inflaton ($\phi$) into a pair of fermions $f$ and $\bar{f}$ via the interaction $y\,\phi\,\bar{f}\,f$, where $y$ is the dimensionless Yukawa coupling. Our analysis focuses on how the non-standard evolution of the Hubble expansion rate and the thermal bath temperature during reheating influence DM production. Additionally, we examine the impact of electroweak symmetry breaking (EWSB), distinguishing between scenarios where DM freeze-in or freeze-out occurs before or after EWSB. We further explore the viable DM parameter space and its compatibility with current and future detection experiments, including XENONnT, LUX-ZEPLIN (LZ), XLZD, and collider searches. Moreover, we incorporate constraints from the Lyman-$\alpha$ bound to ensure consistency with small-scale structure formation.
@@ I Introduction
One compelling alternative is the freeze-in mechanism McDonald (2002); Hall et al. (2010); Elahi et al. (2015); Bernal et al. (2017), where DM interacts so feebly with the SM bath that it never reaches thermal equilibrium. Instead, its relic abundance is generated gradually through the decay and annihilations of SM particles. Due to their extremely weak couplings, such DM candidates are often referred to as feebly interacting massive particles (FIMPs). Another simple way to evade the experimental constraints on WIMP is to consider non-standard cosmological histories, such as the post-inflationary reheating epoch. In recent years, the study of non-standard cosmological histories has gained substantial interest in DM phenomenology Maity and Saha (2018); Garcia et al. (2020, 2021a, 2021b); Giudice et al. (2001); Barman et al. (2022); Bhattiprolu et al. (2022); Harigaya et al. (2014, 2019); Okada and Seto (2021); Ghosh et al. (2022); Haque and Maity (2022); Ahmed et al. (2023); Bernal and Xu (2022); Bernal et al. (2024a, 2019a, 2019b); Bhatia and Mukhopadhyay (2021); Hamdan and Unwin (2018); Arcadi et al. (2024); Barman et al. (2024a); González et al. (2024a, b); Silva-Malpartida et al. (2025); Bernal et al. (2024b); Barman et al. (2024b); Cosme et al. (2024a, b); Haque et al. (2024a, b); Freese et al. (2024); Boddy et al. (2025); Chowdhury and and Show (2025); Addazi and Khlopov (2017); Chowdhury and Hait (2023); Banerjee et al. (2025).
@@
In our work, we investigate the production of Majorana fermionic DM ($\chi$) via the Higgs portal interaction $\bar{\chi}\chi H^{\dagger}H/\Lambda$, where $H$ represents the SM Higgs doublet, and $\Lambda$ serves as the cutoff scale. While some recent studies Lebedev et al. (2024); Arcadi et al. (2024); Ikemoto et al. (2023); Haba et al. (2024); Biswas et al. (2020) have explored the impact of reheating on Majorana DM production, they often assume instantaneous reheating scenarios and treat the reheating temperature as the maximum bath temperature. However, in realistic inflationary models, reheating occurs over a finite duration, where the maximum temperature can be significantly higher than the reheating temperature Giudice et al. (2001), and with a distinct thermal evolution during this phase. Since DM production can depend not only on these characteristic temperatures but also on the thermal history of the Universe. Therefore, our analysis adopts a non-instantaneous reheating scenario to capture these effects more accurately. Furthermore, we consider both freeze-in and freeze-out production mechanisms during the reheating phase. During this phase, the inflaton field ($\phi$) oscillates around the minima of its potential and gradually decays into SM particles, thereby reheating the Universe. We assume that the inflaton oscillates with a monomial potential, $V(\phi)\propto\phi^{2\,n}$, leading to an average equation of state (EoS) given by $w_{\rm\phi}=(n-1)/(n+1)$ Turner (1983). Depending on the choice of $n$ and the inflaton decay width, the energy densities of the inflaton and the radiation bath follow a non-trivial evolution during reheating. This modified cosmological history can significantly alter the DM production process, affecting both its relic abundance and detection prospects.
@@ II Reheating dynamics
After the end of inflation, the inflaton field $(\phi)$ starts to oscillate around the minima of its potential, and transfers its energy into massless fields known as radiation. The process of transferring energy from the inflaton field to the radiation is called reheating. Here, we assume that the inflaton potential during reheating takes the form $V(\phi)\propto\phi^{2\,n}$ with n being an integer, which could originate from the $\alpha$-attractor inflation models Kallosh and Linde (2013); Kallosh et al. (2013)
@@
Here, $M_{\rm pl}(=2.4\times 10^{18}$ GeV) is the reduced Planck mass, $\lambda$ is the potential scale determined from the CMB measurements Garcia et al. (2020), and the parameter $(\alpha,n)$ controls the shape of the potential. The coherently oscillating inflaton field can be decomposed as Shtanov et al. (1995); Ichikawa et al. (2008)
@@
$\omega=m_{\phi}(t)\,\xi\,,~~~\mbox{where}~~\xi=\sqrt{\frac{\pi\,n}{(2\,n-1)}}\frac{\Gamma\left(\frac{1}{2}+\frac{1}{2\,n}\right)}{\Gamma\left(\frac{1}{2\,n}\right)}\,.$ (2)
@@
$m_{\phi}^{\rm end}\simeq\sqrt{\frac{{2n\,\left(2n-1\right)}}{{3\,\alpha}}}\frac{\Lambda^{\frac{2}{n}}}{M_{\rm pl}}\left(\rho_{\phi}^{\rm end}\right)^{\frac{n-1}{2n}}\,.$ (4)
The average energy density of the inflaton ($\rho_{\rm\phi}$) can be defined in terms of the slowly varying amplitude $\phi_{0}$ as $\rho_{\phi}=(1/2)\langle({\dot{\phi}}^{2}+V(\phi))\rangle=V(\phi)|_{\phi_{0}}$. During reheating, the inflaton transfers its energy density to SM radiation energy density $(\rho_{\rm R})$. In order to solve reheating dynamics, the Boltzmann equations for the inflaton and radiation energy density supplemented with the Hubble equation are,
@@
where, dots denote derivatives with respect to cosmic time $t$, and $\Gamma_{\phi}$ is the inflaton decay rate. We consider inflaton decays into a pair of SM fermions $f$ and $\bar{f}$ through the interaction $y\,\phi\,\bar{f}\,f$, where $y$ is the dimensionless Yukawa coupling. Incorporating the inflaton oscillation effect, the effective inflaton decay rate $\Gamma_{\phi}$ can be written as Garcia et al. (2021a); Haque et al. (2023); Silva-Malpartida et al. (2023); Bernal and Xu (2022),
@@
The numerical values of $\sum\nu^{3}\lvert\mathcal{P}_{\nu}\rvert^{2}$ for different values of $w_{\rm\phi}$ are provided in table-1. Incorporating all these parameters, we have numerically solved the coupled Boltzmann equation (Eq. 5) with the appropriate initial conditions: $\rho_{\phi}(A_{\rm end}=1)=\rho_{\phi}^{\rm end}=(3/2)\,V(\phi_{\rm end})\,,\rho_{R}(A_{\rm end}=1)=0$ (see Appendix-A for derivation). For the analytical solution, we assume the inflaton decay width remains negligible compared to the Hubble rate throughout the reheating phase. Under this assumption, the decay term on the right-hand side of Eq. 5a can be neglected, and a straightforward integration yields the following solution,
@@
where $H_{\rm end}$ denotes the Hubble constant at the end of the inflation. In thermal equilibrium, the temperature of the inflaton decay products is given by $\rho_{\rm R}=\epsilon\,T^{4}$, where $\epsilon=(\pi^{2}/30)\,g_{\star}$, with $g_{\star}$ representing the effective number of relativistic degrees of freedom in the thermal bath. For an analytical estimation of the temperature, we assume $g_{\star}\sim\mathcal{O}(100)$, which remains a reasonable approximation for temperatures above $1$ GeV. In our numerical analysis, we solve the following equation along with Eqs. 5, to obtain the correct evolution of the temperature,
@@
$T_{\rm max}$ $=$ $\left[\frac{6\,M_{\rm pl}^{2}\,(1+w_{\phi})m^{end}_{\phi}\,H_{end}}{8\,\pi\,\epsilon\,(5-9w_{\phi})}y_{\rm eff}^{2}\right]^{1/4}\left[\left(\frac{8}{3+9w_{\phi}}\right)^{\frac{3(1+3w_{\phi})}{9w_{\phi}-5}}-\left(\frac{8}{3+9w_{\phi}}\right)^{\frac{8}{9w_{\phi}-5}}\right]^{1/4}$ (12)
@@
To avoid disrupting Big Bang Nucleosynthesis (BBN), the reheating temperature must satisfy $T_{\rm re}\gtrsim T_{\rm BBN}=4$ MeV Kawasaki et al. (2000); Hannestad (2004); Barbieri et al. (2025); de Salas et al. (2015). The upper bound on the reheating temperature is set by the inflationary scale, with $T_{\rm re}\lesssim 10^{15}$ GeV, derived from the constraint on the Hubble parameter during inflation, $H_{\rm I}^{\rm CMB}\leq 2\times 10^{-5}M_{\rm pl}$ Akrami et al. (2020); Ade et al. (2021) (see Appendix-B for derivation). However, the thermal bath temperature can be expressed in terms of $T_{\rm re}$,
${}T(a)\simeq T_{\rm re}\left\{\begin{array}[]{ll}&\left(\frac{A}{A_{\rm re}}\right)^{-\frac{3}{8}(1+3w_{\rm\phi})}~~~\mbox{for}~~A_{\rm max}\leq A\leq A_{\rm re}\,,\\ &\left(\frac{A}{A_{\rm re}}\right)^{-1}~~~\mbox{for}~~A\geq A_{\rm re}\,.\end{array}\right$. (17)
@@
[…]
@@ III.1 DM production before EWSB
$\langle\sigma v\rangle=\frac{g_{\rm\chi}^{2}}{n^{2}_{\rm eq}}\frac{T}{32\pi^{4}}\int^{\infty}_{4m^{2}_{\rm\chi}}\sqrt{s}(s-4m^{2}_{\rm\chi})\sigma(s)K_{1}\left(\frac{\sqrt{s}}{T}\right)ds\,$, (25)
@@
$\langle\sigma v\rangle=\frac{g_{\rm\chi}^{2}}{n_{\text{eq}}^{2}}\cdot\frac{T^{6}}{32\,\pi^{5}\,\Lambda^{2}}\,I_{1}(m_{\chi},T)$ (26)
@@ III.1.1 Freeze-in
$\frac{dN_{\chi}(A)}{dA}=\frac{A^{2}}{H}\langle\sigma v\rangle(n_{\rm eq})^{2}.$ (27)
@@
$N_{\rm{\chi}}(A)=\frac{g_{\rm\chi}^{2}}{24\,(3-7\,w_{\rm\phi})\,\pi^{5}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T_{\rm re}^{6}}{H_{\rm re}}\left[\left(\frac{A}{A_{\text{re}}}\right)^{\frac{3(3-7w_{\phi})}{4}}-\left(\frac{A_{\rm max}}{A_{\text{re}}}\right)^{\frac{3(3-7w_{\phi})}{4}}\right]$ (29)
From the above solution, $3-7w_{\rm\phi}>0$ when $w_{\rm\phi}<3/7$, therefore most of the production occurs around $A\sim A\,(T\sim m_{\rm\chi})$, consequently, the freeze-in temperature $T_{\rm fi}\sim m_{\chi}$ [^1] This type of freeze-in scenario is known as infra-red (IR) freeze-in Bernal et al. (2017); Yaguna (2011); Chu et al. (2012); Blennow et al. (2014); Merle and Totzauer (2015); Shakya (2016); Hessler et al. (2017); König et al. (2016); Duch et al. (2018); Heeba et al. (2018); Lebedev and Toma (2019), where bulk of the DM production occurs at $T\sim m_{\rm\chi}$.. But, on the other hand, $3-7w_{\rm\phi}<0$ when $w_{\rm\phi}>3/7$, most of the production occurs around $A\sim A_{\rm max}$ and this leads to the freeze-in temperature $T_{\rm fi}\sim T_{\rm max}$  [^2] DM production that occurs predominantly at very high temperatures, typically near the maximum temperature of the Universe, is referred to as Ultraviolet (UV) freeze-in Hall et al. (2010); Elahi et al. (2015); McDonald (2016); Chen and Kang (2018); Biswas et al. (2020); Bernal et al. (2019b, 2020a, 2020b); Barman et al. (2020, 2021).. These two scenarios are illustrated in Fig. 3 (dashed lines). In the left plot, for $w_{\rm\phi}=0$, the co-moving number density is frozen at $m_{\chi}/T\sim 1$, and in the right side, for $w_{\rm\phi}=0.5$, the co-moving number density is frozen at the beginning of reheating (at $T\sim T_{\rm max}$).
Figure 3: The evolution of the co-moving number density of DM for freeze-in as a function of the scale factor $A$ (bottom x-axis) and the bath temperature $T$ (top x-axis) for $w_{\rm\phi}=0.0$ (left), $w_{\rm\phi}=0.5$ (right) with $T_{\rm re}=10^{9}$ GeV. Each plot presents three different cases, Left panel: $m_{\rm\chi}(=10^{6}\,\mbox{GeV})<T_{\rm re}$ (solid line, $\Lambda=8.4\times 10^{18}$ GeV), $T_{\rm re}<m_{\rm\chi}(=10^{11}\,\mbox{GeV})<T_{\rm max}$ (dashed line, $\Lambda=3.6\times 10^{16}\,\mbox{GeV}$) and $m_{\rm\chi}(=5.0\times 10^{12}\,\mbox{GeV})>T_{\rm max}$ (dot-dashed, $\Lambda=2.0\times 10^{12}\,\mbox{GeV}$); Right panel: $m_{\rm\chi}(=10^{6}\,\mbox{GeV})<T_{\rm re}$ (solid line, $\Lambda=7.2\times 10^{19}$ GeV), $T_{\rm re}<m_{\rm\chi}(=10^{11}\,\mbox{GeV})<T_{\rm max}$ (dashed line, $\Lambda=2.0\times 10^{22}\,\mbox{GeV}$) and $m_{\rm\chi}(=1.5\times 10^{15}\,\mbox{GeV})>T_{\rm max}$ (dot-dashed, $\Lambda=7.5\times 10^{16}\,\mbox{GeV}$). The vertical red line corresponds to $T\sim T_{\rm fi}$, and the vertical black line corresponds to $T=T_{\rm re}$.
@@
$N_{\rm{\chi}}(A)=\frac{g_{\rm\chi}^{2}\,m_{\rm\chi}}{64\,(1+3\,w_{\rm\phi})\,\pi^{4}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T^{5}_{\text{re}}}{H_{\text{re}}}\left(\frac{T_{\text{re}}}{T_{\text{max}}}\right)^{\frac{7-11w_{\rm\phi}}{1+3w_{\rm\phi}}}e^{-\frac{2\,m_{\rm\chi}}{T_{\rm max}}}$ (30)
@@
$N_{\rm{\chi}}(A)=\frac{g_{\rm\chi}^{2}}{32\,\pi^{5}\,\Lambda^{2}}\frac{A^{3}_{\text{re}}T^{6}_{\text{re}}}{H_{\text{re}}}\left(1-\frac{A_{\text{re}}}{A}\right)$ (31)
Although $m_{\rm\chi}<T_{\rm re}$, the DM production continues up to $T\sim T_{\rm re}$ point, then its production becomes negligible. This is true for $w_{\rm\phi}<3/7$ (see solid line in the left Fig. 3) and the freeze-in temperature $T_{\rm fi}\sim T_{\rm re}$. On the other hand, when $w_{\rm\phi}>3/7$, the DM production is frozen just at the beginning of reheating, which we discussed earlier (see solid line in the right Fig. 3) and the freeze-in temperature $T_{\rm fi}\sim T_{\rm max}$ .
@@
$\Omega_{\rm{\chi}}=\Omega_{R}h^{2}\frac{\sqrt{3}M_{\rm pl}\,g_{\rm\chi}^{2}}{24\,\pi^{5}\epsilon^{\frac{3}{2}}\,T_{0}}\left\{\begin{array}[]{ll}&\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\left(\frac{m_{\rm\chi}}{T_{\rm re}}\right)^{\frac{14\,w_{\phi}-6}{1+3w_{\phi}}}\,,~~~\mbox{for}~~~T_{\rm max}>m_{\rm\chi}>T_{\rm re}\,,w_{\rm\phi}<3/7\\ &\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\,,~~~\mbox{for}~~~m_{\rm\chi}<T_{\rm re}\,,w_{\rm\phi}<3/7\\ &\frac{m_{\chi}T_{\text{re}}}{(3-7w_{\phi})\,\Lambda^{2}}\left(\frac{T_{\rm max}}{T_{\rm re}}\right)^{\frac{(7\,w_{\phi}-3)}{4}}\,,~~~\mbox{for}~~~m_{\rm\chi}<T_{\rm max}\,,w_{\rm\phi}>3/7\\ &\frac{3\pi}{8}\frac{m_{\chi}^{2}}{(1+3w_{\phi})\Lambda^{2}}\left(\frac{T_{\text{re}}}{T_{\text{max}}}\right)^{\frac{7-11w_{\rm\phi}}{1+3w_{\rm\phi}}}e^{-\frac{2\,m_{\rm\chi}}{T_{\rm max}}}\,,~~~\mbox{for}~~m_{\rm\chi}>T_{\rm max}\,.\end{array}\right$. (36)